OpenStudy (anonymous):

find the inverse of the function: f(x)= 3/x^1/3

OpenStudy (anonymous):

\[ y = \frac{3}{x^{1/3}} \]You can start by cubing both sides: \[ y^3 = \frac{3^3}{x^{3/3}} \implies y^3 = \frac{27}{x}\]Now, solve for x :-)